OpenStudy (anonymous):

PLEASE HELP Verify the identity (cotx+tanx)cscxsecx=sec^2x+csc^2x

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

my bet is that this has very little to do with trig and a lot to do with algebra best bet is to rewrite everything in terms of sine and cosine, do the algebra and see if it works out want to try it?

OpenStudy (anonymous):

oh yes, it is algebra for sure

OpenStudy (anonymous):

it will be a lot easier to write if we replace \(\cos(x)\) by \(a\) and \(\sin(x)\) by \(b\) then we can rewrite the left hand side as \[\frac{\frac{a}{b}+\frac{b}{a}}{\frac{1}{ab}}\] is that clear so far?

OpenStudy (anonymous):

let me know if it is or is not, then we can continue only one algebra step to finish this

OpenStudy (anonymous):

im sorry im lost. i dont get anything that deals with this.

OpenStudy (anonymous):

ok we can go slow do you know what \(\cot(x)\) is in terms of \(\sin(x)\) and \(\cos(x)\)? in other words, the basic definition of cotangent?

OpenStudy (anonymous):

"no" is a perfectly good answer, i can tell you if you like

OpenStudy (loser66):

start from the left hand side (cot + tan) csc sec = cot csc sec + tan csc sec ( I just open the bracket) the first term is cot csc sec = \(\dfrac {cos}{sin}*csc *\dfrac{1}{cos}= \dfrac{1}{sin}*csc =csc^2\) the second term is tan csc sec = \(\dfrac{sin}{cos}*\dfrac{1}{sin}*sec=\dfrac{1}{cos}*sec = sec^2\) ------------------------------------------------- add 2 terms, you get the right hand side.